The following table compares the completion percentage and interception percentage of 55 NFL quarterbacks.

Completion Percentage 57 60 62 63 64
Interception Percentage 4.7 4.4 3.5 2.5 2.4

Step 1 of 5 :  

Calculate the sum of squared errors (SSE). Use the values b0=25.3545 and b1=−0.3571 for the calculations. Round your answer to three decimal places.

Step 2 of 6: Calculate the estimated variance of errors, s2e. Round your answer to three decimal places.

Answer =

Step 3 of 6: Calculate the estimated variance of slope, s2b1. Round your answer to three decimal places.

Answer =

Step 4 of 6 : Construct the 90% confidence interval for the slope. Round your answers to three decimal places.

Lower endpoint??

Upper Endpoint??

Step 5 of 6 :  

Construct the 95% confidence interval for the slope. Round your answers to three decimal places.

Lower end point?

upper endpoint?

Step 6 of 6: Construct the 80% confidence interval for the slope. Round your answers to three decimal places.

Lower ?

Upper?
Answer #1
S.No X Y (x-x̅)2 (y-y̅)2 (x-x̅)(y-y̅)
1 57 4.7 17.6400 1.4400 -5.0400
2 60 4.4 1.4400 0.8100 -1.0800
3 62 3.5 0.6400 0.0000 0.0000
4 63 2.5 3.2400 1.0000 -1.8000
5 64 2.4 7.8400 1.2100 -3.0800
Total 306 17.5 30.8000 4.4600 -11.0000
Mean 61.200 3.500 SSX SSY SXY

Step 1 of 5 :  

SSE =Syy-(Sxy)2/Sxx= 0.531

Step 2 of 6:

error Variance σ2 = s2 =SSE/(n-2) = 0.177

Step 3 of 6: Calculate the estimated variance of slope, s2b1.=s2/SSX =0.006

Step 4 of 6 : Construct the 90% confidence interval for the slope:

for 90 % CI value of t= 2.3530
margin of error E=t*std error                            = 0.178
lower confidence bound=estimated slope-margin of error = -0.536
Upper confidence bound=estimated slope+margin of error= -0.179

Step 5 of 6 :  

Construct the 95% confidence interval for the slope :

lower confidence bound=estimated slope-margin of error = -0.598
Upper confidence bound=estimated slope+margin of error= -0.116

Step 6 of 6: Construct the 80% confidence interval for the slope:

lower confidence bound=estimated slope-margin of error = -0.481
Upper confidence bound=estimated slope+margin of error= -0.233